# Stong order $1$ adaptive approximation of jump-diffusion SDEs with discontinuous drift

### v1

@@ 3.3 Convergence result
Now we are able to prove the convergence rate for the jump-adapted quasi-Milstein scheme in the following theorem. In the proof we make use of ideas from [57, Theorem 3.14] and [65, Theorem 1]. We prove the following auxiliary lemma first.
@@
${\mathbb{E}}\Big{[}\sup_{t\in[0,T]}\Big{|}\int_{0}^{s}\widetilde{\sigma}(Z^{(\delta)}_{\underline{u}})d_{\widetilde{\mu}}(Z^{(\delta)}_{\underline{u}})(W_{u}-W_{\underline{u}})\mathop{}\!\mathrm{d}u\Big{|}^{p}\Big{]}^{\frac{1}{p}}\leq c_{9}\delta.$ (3.82)
@@
In the following we prove that $(U_{\delta,\tau_{n}})_{n\in{\mathbb{N}}}$ is a discrete martingale with respect to the filtration $(\widetilde{\mathbb{F}}_{\tau_{n}})_{n\in{\mathbb{N}}}$. By Lemma 3.7 and Lemma 3.9, for all $n\in{\mathbb{N}}$,
@@
Note that $U_{\delta,\tau_{n}}$ is $\widetilde{\mathbb{F}}_{\tau_{n}}$-measurable for all $n\in{\mathbb{N}}$. Hence it remains to prove the martingale property. For this recall that by the proof of Lemma 3.7 we can express $\tau_{n+1}$ as a measurable function $f$ of $\tau_{n}$, $Z^{(\delta)}_{\tau_{n}}$, and $(N_{\tau_{n}+s}-N_{\tau_{n}})_{s\geq 0}$ and calculate for fixed $n\in{\mathbb{N}}$,
@@
${\mathbb{P}}$-a.s. Observe that
@@
Next we estimate the expectation of the $p$-th moment separately. For the first summand we obtain by the discrete Burkholder-Davis-Gundy inequality and Jensen’s inequality that there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that
@@
For the expectation of the second summand of (3.87) we calculate with similar steps as in (3.88) that there exists a constant $\widetilde{c}_{2}\in(0,\infty)$
@@
Combining (3.87), (3.88), and (3.89) we obtain that there exists a constant $c_{9}\in(0,\infty)$ such that
@@
By Hölder’s inequality it is enough to prove the claim for $p\in{\mathbb{N}}\setminus\{1\}$. For all $s\in[0,T]$,
@@
We estimate the first summand by
@@
For all $u\in[0,T]$ we obtain using Remark 3.2,
@@
Using this and Jensen’s inequality we get for the first summand of (3.93) that there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that
@@
For the first summands of (3.95) we get
@@
For the second summand of (3.95) we use Lemma 3.9 to obtain
@@
For the third summand of (3.95) we use Lemma 3.14 to obtain
@@
For the fourth summand of (3.95) we get by Lemma 3.9
@@
For the fifth summand of (3.95) we use Lemma 3.7 and Lemma 3.9 to obtain
@@
Plugging (3.96), (3.97), (3.98), (3.99), and (3.100) into (3.95) we obtain that there exists $\widetilde{c}_{2}\in(0,\infty)$ such that
@@
$\leq\widetilde{c}_{2}\,\int_{0}^{s}{\mathbb{E}}\Big{[}\sup_{v\in[0,u]}\big{|}Z_{u}-Z^{(\delta)}_{u}\big{|}^{p}\Big{]}\mathop{}\!\mathrm{d}u+\widetilde{c}_{2}\delta^{p}.$
For the second summand of (3.93) we apply Lemma 3.15. Hence we obtain that there exists a constant $\widetilde{c}_{3}\in(0,\infty)$ such that
@@
$\leq\widetilde{c}_{3}\,\int_{0}^{s}{\mathbb{E}}\Big{[}\sup_{v\in[0,u]}\big{|}Z_{u}-Z^{(\delta)}_{u}\big{|}^{p}\Big{]}\mathop{}\!\mathrm{d}u+\widetilde{c}_{3}\delta^{p}.$
For estimating the second summand of (3.92), observe that for $u\in(\tau_{n},\tau_{n+1}]$ by Remark 3.2,
@@
Using this, the Burkholder-Davis-Gundy inequality, and Jensen’s inequality we obtain for the second summand of (3.92) that there exist constants $\widetilde{c}_{4},\widetilde{c}_{5}\in(0,\infty)$ such that
@@
Combining (3.104), (3.96), (3.97), (3.98), (3.99), and (3.100) we obtain that there exists a constant $\widetilde{c}_{6}\in(0,\infty)$ such that
@@
$\leq\widetilde{c}_{6}\,\int_{0}^{s}{\mathbb{E}}\Big{[}\sup_{v\in[0,u]}\big{|}Z_{u}-Z^{(\delta)}_{u}\big{|}^{p}\Big{]}\mathop{}\!\mathrm{d}u+\widetilde{c}_{6}\,\delta^{p}.$
For the third summand of (3.92), Lemma 3.6 ensures
@@
Plugging (3.102), (3.105), and (3.106) into (3.92) we obtain that there exists a constant $\widetilde{c}_{7}\in(0,\infty)$ such that
@@
Note that ${\mathbb{E}}\Big{[}\sup_{v\in[0,T]}\big{|}Z_{v}-Z^{(\delta)}_{v}\big{|}^{p}\Big{]}<\infty$ and that $s\mapsto{\mathbb{E}}\Big{[}\sup_{v\in[0,s]}\big{|}Z_{v}-Z^{(\delta)}_{v}\big{|}^{p}\Big{]}$ is a Borel measurable mapping, because it is monotonically increasing. Hence we apply Gronwall’s inequality and obtain that there exists a constant $c_{10}\in(0,\infty)$ such that
@@ 3.4 Cost Analysis
to determine the dependence of the computational cost on $\delta$. The same quantity has, for example, been studied in [44, 65].
@@
$\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{\tau_{k+1}-\tau_{k}}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathop{}\!\mathrm{d}t=\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}.$ (3.111)
@@
$n(Z^{(\delta)}_{T})=\sum_{k\in{\mathbb{N}}}\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{\tau_{k+1}-\tau_{k}}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathop{}\!\mathrm{d}t$ (3.112)
$=\sum_{k\in{\mathbb{N}}}\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{\tau_{k+1}-\tau_{k}}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathds{1}_{\{N_{\tau_{k+1}}-N_{\tau_{k}}=1\}}\mathop{}\!\mathrm{d}t$
$\quad+\sum_{k\in{\mathbb{N}}}\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{\tau_{k+1}-\tau_{k}}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathds{1}_{\{N_{\tau_{k+1}}-N_{\tau_{k}}=0\}}\mathop{}\!\mathrm{d}t$
@@
$\quad+\sum_{k\in{\mathbb{N}}}\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{h^{\delta}(Z^{(\delta)}_{\tau_{k}})}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathds{1}_{\{N_{\tau_{k+1}}-N_{\tau_{k}}=0\}}\mathop{}\!\mathrm{d}t$
$\leq\sum_{k\in{\mathbb{N}}}\mathds{1}_{\{N_{\tau_{k+1}}-N_{\tau_{k}}=1\}}+\sum_{k\in{\mathbb{N}}}\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{h^{\delta}(Z^{(\delta)}_{\tau_{k}})}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathop{}\!\mathrm{d}t$
@@
$\quad+\int_{0}^{T}\frac{1}{h^{\delta}(Z^{(\delta)}_{\underline{t}})}\mathds{1}_{\Theta^{\varepsilon_{1}^{\delta}}\setminus\Theta^{\varepsilon_{2}^{\delta}}}(Z^{(\delta)}_{\underline{t}})\mathop{}\!\mathrm{d}t+\int_{0}^{T}\frac{1}{h^{\delta}(Z^{(\delta)}_{\underline{t}})}\mathds{1}_{\Theta^{\varepsilon_{2}^{\delta}}}(Z^{(\delta)}_{\underline{t}})\mathop{}\!\mathrm{d}t$
Now we estimate the expectation of each summand separately. We calculate
@@
Further, using similar calculations to (3.80) we get that there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that
@@
It remains to prove that there exists a constant $\widetilde{c}_{2}\in(0,\infty)$ such that for all $\delta\in[0,\delta_{0})$
@@
then (3.112), (3.113), (3.114), and (3.115) imply the claim. To prove (3.115) we define the set
@@
Since the distance function $d(\cdot,\Theta)\colon{\mathbb{R}}\to[0,\infty)$ is Lipschitz continuous with Lipschitz constant 1, we get
@@
[…]
Now we define $\bar{\varepsilon}^{\delta}=\delta^{3/4}\log^{3}(\delta^{-1})$ and note that $\varepsilon_{2}^{\delta}\leq\overline{\varepsilon}^{\delta}\leq\varepsilon_{1}^{\delta}$ for all $\delta\in(0,\delta_{0}]$. This implies
@@
[…]
@@
Applying Lemma 3.11 choosing $f=(\max\{\overline{\varepsilon}^{\delta},\cdot\})^{-2}$ and $\gamma=1/2$ resp. $f=(\max\{\tfrac{1}{2}\varepsilon_{2}^{\delta},\cdot\})^{-2}$ and $\gamma=1/6$ we obtain that there exists a constant $\widetilde{c}_{3}\in(0,\infty)$ and a constant $\widetilde{c}_{4}\in(0,\infty)$,
@@
Next we use that $\delta^{2}\log^{4}(\delta^{-1})\leq h^{\delta}$ and Lemma 3.12 (ii) with $\alpha=1/2$ and $q=2$ to obtain that there exists a constant $\widetilde{c}_{5}\in(0,\infty)$

### v2

@@ 3.3 Convergence result
Now we are able to prove the convergence rate for the jump-adapted quasi-Milstein scheme in the next theorem. The proof of the following two results follows the general strategy introduced in [38, Proof of Theorem 3], which was extended in [65, Section 5.4] to adaptive step-size functions and in [57, Proof of Theorem 3.14] to jump-adapted schemes. We make use of the general strategy as well as both extensions to prove for SDE (3.1) and the approximation scheme (3.11) the same convergence result as in [65]. We prove the following auxiliary lemma first.
@@
${\mathbb{E}}\Big[\sup_{s\in[0,T]}\Big|\int_{0}^{s}\widetilde{\sigma}(Z^{(\delta)}_{\underline{u}})d_{\widetilde{\mu}}(Z^{(\delta)}_{\underline{u}})(W_{u}-W_{\underline{u}})\mathop{}\!\mathrm{d}u\Big|^{p}\Big]^{\frac{1}{p}}\leq c_{9}\delta.$ (3.83)
@@
In the following we prove that $(U_{\delta,\tau_{n}})_{n\in{\mathbb{N}}}$ is a discrete martingale with respect to the filtration $(\widetilde{\mathbb{F}}_{\tau_{n}})_{n\in{\mathbb{N}}}$. By Lemma 3.7 and Lemma 3.9, for all $n\in{\mathbb{N}}$ and all $\delta\in(0,\delta_{0}]$,
@@
Note that $U_{\delta,\tau_{n}}$ is $\widetilde{\mathbb{F}}_{\tau_{n}}$-measurable for all $n\in{\mathbb{N}}$. Hence it remains to prove the martingale property. For this recall that by the proof of Lemma 3.7 we can express $\tau_{n+1}$ as a measurable function $f$ of $\tau_{n}$, $Z^{(\delta)}_{\tau_{n}}$, and $(N_{\tau_{n}+s}-N_{\tau_{n}})_{s\geq 0}$ and calculate for all $n\in{\mathbb{N}}$ and all $\delta\in(0,\delta_{0}]$,
@@
$\mathbb{P}$}-almost surely. Observe that for all $\delta\in(0,\delta_{0}]$,
@@
Next we estimate the expectation of the $p$-th moment separately. For the first summand we obtain by the discrete Burkholder-Davis-Gundy inequality and Jensen’s inequality that there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$,
@@
For the expectation of the second summand of (3.88) we calculate with similar steps as in (3.89) that there exists a constant $\widetilde{c}_{2}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$,
@@
Combining (3.88), (3.89), and (3.90) we obtain that there exists a constant $c_{9}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$,
@@
By Hölder’s inequality it is enough to prove the claim for $p\in{\mathbb{N}}\setminus\{1\}$. For all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
We estimate the first summand for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$ by
@@
For all $u\in[0,T]$ and all $\delta\in(0,\delta_{0}]$ we obtain using Remark 3.2,
@@
Using this and Jensen’s inequality, we get for the first summand of (3.94) that there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
For the first summands of (3.96) we get for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
For the second summand of (3.96) we use Lemma 3.9 to obtain for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
For the third summand of (3.96) we use Lemma 3.14 to obtain for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
For the fourth summand of (3.96) we get by Lemma 3.9 for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
For the fifth summand of (3.96) we use Lemma 3.7 and Lemma 3.9 to obtain for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
Plugging (3.97), (3.98), (3.99), (3.100), and (3.101) into (3.96) we obtain that there exists $\widetilde{c}_{2}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
$\leq\widetilde{c}_{2}\,\int_{0}^{s}{\mathbb{E}}\Big[\sup_{v\in[0,u]}\big|Z_{v}-Z^{(\delta)}_{v}\big|^{p}\Big]\mathop{}\!\mathrm{d}u+\widetilde{c}_{2}\delta^{p}.$
For the second summand of (3.94) we apply Lemma 3.15. Hence we obtain that there exists a constant $\widetilde{c}_{3}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
$\leq\widetilde{c}_{3}\,\int_{0}^{s}{\mathbb{E}}\Big[\sup_{v\in[0,u]}\big|Z_{v}-Z^{(\delta)}_{v}\big|^{p}\Big]\mathop{}\!\mathrm{d}u+\widetilde{c}_{3}\delta^{p}.$
For estimating the second summand of (3.93), observe that for all $u\in(\tau_{n},\tau_{n+1}]$ and all $\delta\in(0,\delta_{0}]$ by Remark 3.2,
@@
Using this, the Burkholder-Davis-Gundy inequality, and Jensen’s inequality we obtain for the second summand of (3.93) that there exist constants $\widetilde{c}_{4},\widetilde{c}_{5}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
Combining (3.105), (3.97), (3.98), (3.99), (3.100), and (3.101) we obtain that there exists a constant $\widetilde{c}_{6}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
$\leq\widetilde{c}_{6}\,\int_{0}^{s}{\mathbb{E}}\Big[\sup_{v\in[0,u]}\big|Z_{v}-Z^{(\delta)}_{v}\big|^{p}\Big]\mathop{}\!\mathrm{d}u+\widetilde{c}_{6}\,\delta^{p}.$
For the third summand of (3.93), Lemma 3.6 ensures for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
Plugging (3.103), (3.106), and (3.107) into (3.93) we obtain that there exists a constant $\widetilde{c}_{7}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $s\in[0,T]$,
@@
Note that ${\mathbb{E}}\Big[\sup_{v\in[0,T]}\big|Z_{v}-Z^{(\delta)}_{v}\big|^{p}\Big]<\infty$ and that $s\mapsto{\mathbb{E}}\Big[\sup_{v\in[0,s]}\big|Z_{v}-Z^{(\delta)}_{v}\big|^{p}\Big]$ is a Borel measurable mapping, because it is monotonically increasing. Hence we apply Gronwall’s inequality and obtain that there exists a constant $c_{10}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$,
@@ 3.4 Cost Analysis
to determine the dependence of the computational cost on $\delta$. The same quantity has, for example, been studied in [44, 65]. The result and proof of the following theorem are closely based on [65, Section 5.5].
@@
$\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{\tau_{k+1}-\tau_{k}+\mathds{1}_{\{\tau_{k+1}=\tau_{k}\}}}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathop{}\!\mathrm{d}t=\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}.$ (3.112)
@@
$n(Z^{(\delta)}_{T})=\sum_{k\in{\mathbb{N}}}\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{\tau_{k+1}-\tau_{k}+\mathds{1}_{\{\tau_{k+1}=\tau_{k}\}}}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathop{}\!\mathrm{d}t$ (3.113)
$=\sum_{k\in{\mathbb{N}}}\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{\tau_{k+1}-\tau_{k}+\mathds{1}_{\{\tau_{k+1}=\tau_{k}\}}}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathds{1}_{\{N_{\tau_{k+1}}-N_{\tau_{k}}=1\}}\mathop{}\!\mathrm{d}t$
$\quad+\sum_{k\in{\mathbb{N}}}\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{\tau_{k+1}-\tau_{k}+\mathds{1}_{\{\tau_{k+1}=\tau_{k}\}}}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathds{1}_{\{N_{\tau_{k+1}}-N_{\tau_{k}}=0\}}\mathop{}\!\mathrm{d}t$
@@
$\quad+\sum_{k\in{\mathbb{N}}}\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{h^{\delta}(Z^{(\delta)}_{\tau_{k}})}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathds{1}_{\{N_{\tau_{k+1}}-N_{\tau_{k}}=0\}}\mathop{}\!\mathrm{d}t.$
This implies for all $\delta\in[0,\delta_{0})$,
$n(Z^{(\delta)}_{T})\leq\sum_{k\in{\mathbb{N}}}\mathds{1}_{\{N_{\tau_{k+1}}-N_{\tau_{k}}=1\}}+\sum_{k\in{\mathbb{N}}}\int_{\tau_{k}}^{\tau_{k+1}}\frac{1}{h^{\delta}(Z^{(\delta)}_{\tau_{k}})}\mathds{1}_{\{\tau_{k+1}\neq\tau_{k}\}}\mathop{}\!\mathrm{d}t$ (3.114)
@@
$\quad+\int_{0}^{T}\frac{1}{h^{\delta}(Z^{(\delta)}_{\underline{t}})}\mathds{1}_{\Theta^{\varepsilon_{1}^{\delta}}\setminus\Theta^{\varepsilon_{2}^{\delta}}}(Z^{(\delta)}_{\underline{t}})\mathop{}\!\mathrm{d}t+\int_{0}^{T}\frac{1}{h^{\delta}(Z^{(\delta)}_{\underline{t}})}\mathds{1}_{\Theta^{\varepsilon_{2}^{\delta}}}(Z^{(\delta)}_{\underline{t}})\mathop{}\!\mathrm{d}t.$
Now we estimate the expectation of each summand separately. We calculate for all $\delta\in[0,\delta_{0})$,
@@
Further, using similar calculations to (3.81) we get that there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that for all $\delta\in[0,\delta_{0})$,
@@
It remains to prove that there exists a constant $\widetilde{c}_{2}\in(0,\infty)$ such that for all $\delta\in[0,\delta_{0})$,
@@
then (3.114), (3.115), (3.116), and (3.117) imply the claim. To prove (3.117) we define for all $\delta\in[0,\delta_{0})$ and all $t\in[0,T]$ the set
@@
Since the distance function $d(\cdot,\Theta)\colon{\mathbb{R}}\to[0,\infty)$ is Lipschitz continuous with Lipschitz constant 1, we get for all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
$\leq\frac{9}{4}\log^{4}(\delta^{-1}){\mathbb{E}}\Big[\int_{0}^{T}\frac{1}{d(Z^{(\delta)}_{t},\Theta)^{2}}\mathds{1}_{\Theta^{\frac{3}{2}\varepsilon_{1}^{\delta}}\setminus\Theta^{\frac{1}{2}\varepsilon_{2}^{\delta}}}(Z^{(\delta)}_{t})\,dt\Bigr].$
Now we define $\bar{\varepsilon}^{\delta}=\delta^{3/4}\log^{3}(\delta^{-1})$ and note that $\varepsilon_{2}^{\delta}\leq\overline{\varepsilon}^{\delta}\leq\varepsilon_{1}^{\delta}$ for all $\delta\in(0,\delta_{0}]$. This implies for all $\delta\in(0,\delta_{0}]$,
@@
$\leq\frac{9}{4}\log^{4}(\delta^{-1})\Big({\mathbb{E}}\Big[\int_{0}^{T}\frac{1}{d(Z^{(\delta)}_{t},\Theta)^{2}}\mathds{1}_{\Theta^{\frac{3}{2}\varepsilon_{1}^{\delta}}\setminus\Theta^{\bar{\varepsilon}^{\delta}}}(Z^{(\delta)}_{t})\,dt\Bigr]+{\mathbb{E}}\Big[\int_{0}^{T}\frac{1}{d(Z^{(\delta)}_{t},\Theta)^{2}}\mathds{1}_{\Theta^{\bar{\varepsilon}^{\delta}}\setminus\Theta^{\frac{1}{2}\varepsilon_{2}^{\delta}}}(Z^{(\delta)}_{t})\,dt\Bigr]\Big)$
@@
Applying Lemma 3.11 choosing $f=(\max\{\overline{\varepsilon}^{\delta},\cdot\})^{-2}$ and $\gamma=1/2$ resp. $f=(\max\{\tfrac{1}{2}\varepsilon_{2}^{\delta},\cdot\})^{-2}$ and $\gamma=1/6$ we obtain that there exist constants $\widetilde{c}_{3},\widetilde{c}_{4}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$,
@@
Next we use that $\delta^{2}\log^{4}(\delta^{-1})\leq h^{\delta}$ and Lemma 3.12 (ii) with $\alpha=1/2$ and $q=2$ to obtain that there exists a constant $\widetilde{c}_{5}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$,